The average age of a father and his five children is 21. If the ages of the children are 10,12,15,17 and 20 years, how old is the father?
Answer by Theo(13342)

average ages of 6 people is 21.
this means the total age is 6 * 21 = 126
the ages of his children are 10, 12, 15, 17, and 20.
the total of those 5 ages is 74.
126 - 74 = 52
the father is 52 years old.
total ages are now 74 + 52 = 126
126 / 6 = 21
average age of the 6 people is 21.
your solution is that the father's age is 52.